how to solve this
substitution method
5x-3y+3=0
6x-5y-2=0

Answer by stanbon(75887)

solve this
substitution method
5x-3y+3=0
6x-5y-2=0
------------------
Solve for "x" in the top equation::
x = (3/5)(y+1)
------
Substitute for "x" and solve for "y":
6x - 5y - 2 = 0
(18/5)(y+1)-5y-2 = 0
-----------
Multiply thru by 5 to get:
18y+18 - 25y - 10 = 0
-7y = -8
y = 8/7
--------
Solve for "x":
x = (3/5)(y+1)
x = (3/5)(15/7)
x = 45/35
x = 9/7
=================
Cheers,
Stan H.
-----------------